What is the height of an equilateral triangle with sides of 10 in long

Answer by edjones(8007)

The altitude divides the triangle into 2 right triangles, each with a hypotenuse of 10" and a base of 5".
a^2+b^2=c^2
a^2+5^2=10^2
a^2+25=100
a^2=75
a=sqrt(75)
=sqrt(25*3)
=5sqrt(3)" the height
.
Ed